The amount of annual interest earned by $7,000 invested at a certain rate is $210 less than $11,000 would earn at a 1% lower rate. At what rate is the $7,000 invested?
Answer by stanbon(75887)

The amount of annual interest earned by $7,000 invested at a certain rate is $210 less than $11,000 would earn at a 1% lower rate. At what rate is the $7,000 invested?
-------------------
Equation::
7000*x = 0.01*11000 - 210
---------------------------
7000x = 110 - 210
---
x = -100/7000
Note: The rate of investment cannot be negative.
Please check the wording of your problem or your post.
========
Cheers,
Stan H.
----------